# Lense-Thirring precession of accretion flows: relativistic versus classical precession

### v1

@@ 5 Effects of spin
As elaborated in works of Gondek-Rosińska et al. (2014), Tsang & Pappas (2016) and Urbancová et al. (2019), for rotating oblate stars, the geodesic vertical epicyclic frequency exceeds the Keplerian frequency for sufficiently high values of $j$ and $q$. In this situation, as illustrated in Figures 1c and 2, which are drawn for $j=0.35$, the LT precession changes orientation at a radius close to the ISCO. Its frequency values are then negative below this radius and positive above it, while $\nu_{\mathrm{LT}}$ again vanishes at infinity. Figure 2 further illustrates the presence of a local radial maximum of the precession frequency. This maximum may be present even when the LT precession frequency only vanishes at infinity and not at radii close to the ISCO.
@@
In Figure 3a, we show several examples of the behavior of $\nu_{\mathrm{LT}}(j,r=r_{\mathrm{ISCO}})$ for realistic values of $\tilde{q}$. The displayed frequencies are relevant for both the inner edge of a thin accretion disc and a slender torus, which also represents a torus of maximum thickness at this radius. Inspecting the Figure, we can see that the precession frequency clearly exhibits a distinct maximum according to relation (14).
@@ 5.1 Situation above ISCO
Situations at radii $r_{\mathrm{3:1}}$, $r_{\mathrm{2:1}}$ and $r_{\mathrm{3:2}}$, that are qualitatively similar to $r_{\mathrm{ISCO}}$, are illustrated within panels b, c and d of Figure 3. The test particle motion (and slender tori) frequencies are shown using full lines, while the non-slender torus cases are indicated by shaded areas corresponding to tori of all thicknesses allowable according to the results of Matuszková et al. (2024b). While at $r_{\mathrm{3:1}}$ almost no quantitative difference exists between the test particle and fluid motion frequencies, at $r_{\mathrm{2:1}}$, the spread of frequencies given by the freedom in the torus thickness is clearly visible. Nevertheless, the behaviour of the fluid is qualitatively similar to the particle motion. At higher radii, represented by $r_{\mathrm{3:2}}$, the differences between the individual values of $\tilde{q}$ are smaller but still apparent. In each panel of Figure 3, we also include the frequencies calculated in the LT spacetime. The frequency values predicted by monotonic functions associated with this spacetime are substantially higher than the maxima associated with the HT spacetime.
Figure 4: Maximum and average value of the LT frequency. Left panel: The maximum of the frequency, $\nu_{\mathrm{MAX}}$, as a function of the angular momentum of the star. Middle panel: The frequency value averaged from $r_{\mathrm{ISCO}}$ to $r_{\mathrm{2:1}}$ as a function of the star’s angular momentum. Right panel: The frequency averaged from $r_{\mathrm{ISCO}}$ to $r_{\mathrm{3:2}}$ as a function of the star’s angular momentum. As in Figure 3, the red lines correspond to the LT metric while the lines in shades of blue correspond to the HT metric and varying $q$. The colour-shaded regions then indicate frequencies relevant to non-geodetic flows and cover the full range of allowable torus thicknesses.
5.2 Maximal and averaged frequencies
@@
Finally, similar behaviour is also exhibited by the geodesic frequency averaged along the innermost parts of the disc, which is important when the observed LF QPOs are associated with fluid flows precessing as a rigid body (e.g. Ingram & Done, 2011). Two examples of such frequencies are illustrated for the radii from ISCO to $r_{2:1}$ (middle panel of Figure 4) and from ISCO to $r_{3:2}$ (right panel of Figure 4). Within these panels, we also indicate the ranges relevant to the frequencies averaged in the same way but for the tori.
@@ 6 Surface
Figure 5: The maximum LT precession frequency as a function of the angular momentum of the star based on specific EoS and a fixed gravitational mass, $M$ (top panels), or a fixed non-rotating mass, $M_{0}$ (bottom panels). Left panels: XMLSLZ(NL3) EoS. Middle panels: VGBCMR(D1M) EoS. Right panels: MIT bag EoS. The coloured solid lines show the values of $\nu_{\mathrm{MAX}}$ when the star’s surface is below the ISCO, while the coloured dashed lines show the values of $\nu_{\mathrm{MAX}}$ when the surface is above the ISCO. The circles on the individual lines enable us to determine the radius at which the maximum displayed frequency is reached. For values of $j$ below that indicated by the black filled circle (or when there is no circle), the frequency corresponds to the edge of the relevant radial interval: the ISCO (solid lines) or the NS surface (dotted lines). For higher values of $j$, below the value indicated by the open circle, the frequency corresponds to the maximum above the edge. Above this value, it again corresponds to the ISCO for the solid lines and the NS surface for the dotted lines. The grey lines then indicate the value of the corresponding local geodesic maximum, $\nu_{\mathrm{max}}^{0}$.
@@ 7 Discussion and conclusions
The works of Pappas (2012), Kluźniak & Rosińska (2013, 2014), Gondek-Rosińska et al. (2014), Pappas (2015), Tsang & Pappas (2016), and Matuszková et al. (2024a, b) have revealed the importance of the interplay between relativistic effects and effects already present in the Newtonian physics of oblate bodies. In the case of LT precession, this interplay takes place between the contributions of relativistic and classical precession, which have opposite directions, and this is crucial for our study utilising the HT spacetimes.

### v2

@@ 5 Effects of spin
As elaborated in works of Gondek-Rosińska et al. (2014), Tsang and Pappas (2016) and Urbancová et al. (2019), for rotating oblate stars, the geodesic vertical epicyclic frequency exceeds the Keplerian frequency for sufficiently high values of $j$ and $q$. In this situation, as illustrated in Figures 1c and 2, which are drawn for $j=0.35$, the LT precession changes orientation at a radius close to the ISCO. Its frequency values are then negative below this radius and positive above it, while $\nu_{\mathrm{LT}}$ again vanishes at infinity. Figure 2 further illustrates the presence of a local radial maximum of the precession frequency. This maximum may be present even when the LT precession frequency only vanishes at infinity and not at radii close to the ISCO.
@@
In Figure 3a, we show several examples of the behaviour of $\nu_{\mathrm{LT}}(j,r=r_{\mathrm{ISCO}})$ for realistic values of $\tilde{q}$. The displayed frequencies are relevant for the inner edge of a thin accretion disc and for a slender torus, which also represents a torus of maximum thickness at this radius. Inspecting the Figure, we can see that the precession frequency clearly exhibits a distinct maximum according to relation (14).
@@ 5.1 Situation above ISCO
Situations at radii $r_{\mathrm{3:1}}$, $r_{\mathrm{2:1}}$, and $r_{\mathrm{3:2}}$, that are qualitatively similar to $r_{\mathrm{ISCO}}$, are illustrated within panels b, c, and d of Figure 3. The test particle motion (and slender tori) frequencies are shown using full lines, while the non-slender torus cases are indicated by shaded areas corresponding to tori of all thicknesses allowable according to the results of Matuszková et al. (2024b). While at $r_{\mathrm{3:1}}$ almost no quantitative difference exists between the test particle and fluid motion frequencies, at $r_{\mathrm{2:1}}$, the spread of frequencies given by the freedom in the torus thickness is clearly visible. Nevertheless, the behaviour of the fluid is qualitatively similar to the particle motion. At larger radii, represented by $r_{\mathrm{3:2}}$, the differences between the individual values of $\tilde{q}$ are smaller but still apparent. In each panel of Figure 3, we also include the frequencies calculated in the LT spacetime. The frequency values predicted by monotonic functions associated with this spacetime are substantially higher than the maxima associated with the HT spacetime.
Figure 4: Maximum and average value of the LT frequency. Left panel: Maximum of the frequency, $\nu_{\mathrm{MAX}}$, as a function of the angular momentum of the star. Middle panel: Frequency value averaged from $r_{\mathrm{ISCO}}$ to $r_{\mathrm{2:1}}$ as a function of the star’s angular momentum. Right panel: Frequency averaged from $r_{\mathrm{ISCO}}$ to $r_{\mathrm{3:2}}$ as a function of the star’s angular momentum. As in Figure 3, the red lines correspond to the LT metric, while the lines in shades of blue correspond to the HT metric and varying $q$. The colour-shaded regions then indicate the frequencies relevant to non-geodetic flows and cover the full range of allowable torus thicknesses.
5.2 Maximum and averaged frequencies
@@
Finally, similar behaviour is also exhibited by the geodesic frequency averaged along the innermost parts of the disc, which is important when the observed LF QPOs are associated with fluid flows precessing as a rigid body (e.g. Ingram and Done, 2011). Two examples of such frequencies are illustrated for the radii from ISCO to $r_{2:1}$ (middle panel of Figure 4) and from ISCO to $r_{3:2}$ (right panel of Figure 4). Within these panels, we also indicate the ranges relevant to the frequencies averaged in the same way but for the tori.
@@ 6 Surface
Figure 5: Maximum LT precession frequency as a function of angular momentum of the star based on specific EoS and a fixed gravitational mass, $M$ (top panels), or a fixed non-rotating mass, $M_{0}$ (bottom panels). Left panels: XMLSLZ(NL3) EoS. Middle panels: VGBCMR(D1M) EoS. Right panels: MIT bag EoS. The coloured solid lines show the values of $\nu_{\mathrm{MAX}}$ when the star’s surface is below the ISCO, while the coloured dashed lines show the values of $\nu_{\mathrm{MAX}}$ when the surface is above the ISCO. The circles on the individual lines enable us to determine the radius at which the maximum displayed frequency is reached. For values of $j$ below that indicated by the black filled circle (or when there is no circle), the frequency corresponds to the edge of the relevant radial interval: the ISCO (solid lines) or the NS surface (dotted lines). For higher values of $j$, below the value indicated by the open circle, the frequency corresponds to the maximum above the edge. Above this value, it again corresponds to the ISCO for the solid lines and the NS surface for the dotted lines. The grey lines then indicate the value of the corresponding local geodesic maximum, $\nu_{\mathrm{max}}^{0}$.
@@ 7 Discussion and conclusions
The works of Pappas (2012), Kluźniak and Rosińska (2013, 2014), Gondek-Rosińska et al. (2014), Pappas (2015), Tsang and Pappas (2016), and Matuszková et al. (2024a, b) have revealed the importance of the interplay between relativistic effects and effects already present in the Newtonian physics of oblate bodies. In the case of LT precession, this interplay takes place between the contributions of relativistic and classical precession, which have opposite directions, and this is crucial for our study utilising the HT spacetimes.